10. Rihanna, Inc. sells watches for $100 per watch. The variable expenses are $40 per unit, and the fixed

expenses total $35,000 per period. By how much will net operating income change if watch sales are expected

to increase by $10,000?

A. $4,000

B. $3,000

C. $24,000

D. $6,000

11. In the above question, what is the breakeven point for Rihanna, Inc. in sales dollars?

A. $22,558

B. $43,555

C. $45,000

D. $58,333
Answer #1

Contribution margin ratio = (100-40)/100 = 60%

10) Increase net income by (10000*60%) = 6000

So answer is d) $6000

11) Break even point = 35000/.60 = 58333

So answer is d) $58333